The Gamer Company is a video game production company that specializes in educational video games for kids. The company’s R&D department is always looking for great ideas for new games. On average, the R&D department generates about 20 new ideas a week. To go from idea to approved product, the idea must pass through the following stages: paper screening (a 1-page document describing the idea and giving a rough sketch of the design), prototype development, testing, and a focus group. At the end of each stage, successful ideas enter the next stage. All other ideas are dropped. The following chart depicts this process, and the probability of succeeding at each stage. Pr(Pass) =0.60 Pr(Pass) =0.50 Pr(Pass) =0.35 Pr(Pass) =0.75 Paper Screening→Prototype,design &development→Testing→Focus Group → Final Decision→ Fail ↓ Fail↓ Fail↓ Fail↓ The paper screening for each idea takes 2 hours of a staff member’s time. After that, there is a stage of designing and producing a prototype. A designer spends 4 hour designing the game in a computer-aided-design (CAD) package. The actual creation of the mock-up is outsourced to one of many suppliers with essentially limitless capacity. It takes 4 days to get the prototype programmed, and multiple prototypes can be created simultaneously. A staff member of the testing team needs 1 days to test an idea. Running the focus group takes 2 hours of a staff member’s time per idea, and only one game is tested in each focus group. Finally, the management team meets for 3 hours per idea to decide if the game should go into production. Available working hours for each staff member are 8 hours per day, 5 days a week. The current staffing plan is as follows: A. Paper screening: 3 staff members. B. Design and Production: 4 staff members. C. Testing: 5 staff members. D. Focus Group: 1 staff member. E. Finally Decision: 1 management team. What is the implied utilization of testing team in percentage?
Answer #1

Average no. of ideas per week =20

Ideas approved per week = 20*0.6*0.5*0.35*0.75 = 1.575 ideas

Maximum time taking stage is Testing therefore it is the bottleneck according to current staffing plan

Capacity per hour = no. of staff/no. of hour

Capacity per week = Capacity per hour*5*8

Demand is lowest of all capacity= 13.3

Paper screening

Capacity= 3/2=1.5 idea/ hour

= 60 idea per week

Demand =13.3

Utilization=0.22

Design & production

Capacity= 4/4=1 idea/ hour

= 40 idea per week

Demand =13.3

Utilization=0.33

Testing

Capacity= 6/16=0.38 idea/ hour

= 15 idea per week

Demand =13.3

Utilization=0.88

Focus group

Capacity= 1/2=0.5 idea/ hour

= 20 idea per week

Demand =20

Utilization=0.66

Final decision

Capacity= 1/3=0.33 idea/ hour

= 13.3 idea per week

Demand =13.3

Utilization=1

Therefore 13.3 ideas per week